OpenStudy (emac):

Implicit Differentiation dy/dx x^2y-x-5y-9=0

OpenStudy (kittiwitti1):

\[\frac{dy}{dx}x^{2}y-x-5y-9=0\]this?

OpenStudy (emac):

yes!

OpenStudy (kittiwitti1):

okay so first off we split everything up in terms of differentiation:\[\frac{dy}{dx}(x^{2}y)-\frac{dy}{dx}(x)-\frac{dy}{dx}(5y)-\frac{dy}{dx}(9)=\frac{dy}{dx}(0)\]

OpenStudy (kittiwitti1):

(I DON'T KNOW IF THIS IS CORRECT)\[2x\times y'-1-5(1)-9=0\]\[2x\times y'-15=0\]\[2x\times y'=15\]\[y'=\frac{15}{2x}\]

Nnesha (nnesha):

x^2y product rule

OpenStudy (kittiwitti1):

whoops

OpenStudy (kittiwitti1):

okay implicit differentiation of \(x^{2}y\) is\[2xy+x^{2}y'\]

OpenStudy (kittiwitti1):

so then you have the full implicit differentiation...\[2xy+x^{2}y'-1-5(1)-9=0\]\[2xy+x^{2}y'-15=0\]\[x^{2}y'=-2xy+15...\]\[y=\frac{-2xy+15}{x^{2}}\]

OpenStudy (kittiwitti1):

did I get it right? lol

Nnesha (nnesha):

the derivative of constant should be 0

OpenStudy (emac):

oh gotcha, thank you!!

OpenStudy (kittiwitti1):

oh... my bad.

OpenStudy (kittiwitti1):

\[2xy+x^{2}y'-1-5(1)-0=0\]\[2xy+x^{2}y'=6\]

Nnesha (nnesha):

\[\large\rm \color{Red}{x^2y} \color{blue}{-x}\color{orange}{-5y}-9=0\] \[\large\rm \color{Red}{ x^2y= 2x \frac{dx}{dx}(y)+ x^2 (1\frac{dy}{dx} )} \] \[\color{blue}{-x = -1 \frac{dx}{dx}}\] \[ \color{orange}{-5y = -5 \frac{dy}{dx}}\] \[\large\rm \color{Red}{ 2x \frac{dx}{dx}(y)+ x^2 (1\frac{dy}{dx} )} \color{blue}{ -1 \frac{dx}{dx}} \color{orange}{ -5 \frac{dy}{dx}} -0=0\] dy/dx=taking the derivative of y with respect to x dx/dx= derivative of x with respect to x

Nnesha (nnesha):

dx/dx=1 so \[\large\rm 2xy+x^2 \frac{dy}{dx}-1 -5\frac{dy}{dx}=0\] solve for `dy/dx`

Nnesha (nnesha):

@emac makes sense ?? familiar with the product rule ??